# Infinity-MM: Scaling Multimodal Performance with Large-Scale and High-Quality Instruction Data

### v1

@@ Abstract
Vision-Language Models (VLMs) have recently made significant progress, but the limited scale and quality of open-source instruction data hinder their performance compared to closed-source models. In this work, we address this limitation by introducing Infinity-MM, a large-scale multimodal instruction dataset with 40 million samples, enhanced through rigorous quality filtering and deduplication. We also propose a synthetic instruction generation method based on open-source VLMs, using detailed image annotations and diverse question generation. Using this data, we trained a 2-billion-parameter VLM, Aquila-VL-2B, achieving state-of-the-art (SOTA) performance for models of similar scale. This demonstrates that expanding instruction data and generating synthetic data can significantly improve the performance of open-source models.
Infinity-MM: Scaling Multimodal Performance with Large-Scale
and High-Quality Instruction Data
Shuhao Gu1thanks: Core contributors with equal contributions., Jialing Zhang1,2*, Siyuan Zhou1,3*, Kevin Yu1,4*, Zhaohu Xing1,5, Liangdong Wang1, Zhou Cao1, Jintao Jia1,4, Zhuoyi Zhang1,4, Yixuan Wang1,4, Zhenchong Hu1,4, Bo-Wen Zhang1, Jijie Li1, Dong Liang1, Yingli Zhao1, Yulong Ao1, Yaoqi Liu4, Fangxiang Feng3, Guang Liu1thanks: Project Lead, liuguang@baai.ac.cn 1BAAI, 2BJTU, 3BUPT, 4ICT/CAS, 5HKUST(GZ)
@@ 1 Introduction
Recently, Vision-Language Models (VLMs) Li et al. (2023); Liu et al. (2024b); Dai et al. (2023); Zhu et al. (2024); Bai et al. (2023b); Wang et al. (2023b); Xiao et al. (2024); OpenAI (2024); Yao et al. (2024); Wang et al. (2024a); Chen et al. (2024b); Li et al. (2024a) have made significant progresses, drawing increasing attention. With the ongoing advancements in foundational language models, multimodal architectures, multimodal training data, and evaluation benchmarks, the capabilities of multimodal models have greatly improved. Among these developments, the expansion of training data scale, the enhancement of data quality, and the optimization of training strategies have emerged as key factors in boosting model performance Liu et al. (2024b, 2023a); Tong et al. (2024); Li et al. (2024a, b). Currently, the two primary methods for data acquisition are manual data collection and annotation, as well as using models to synthesize instructions.
@@
Many works have focused on exploring more effective ways to generate and utilize training data. For instance, Liu et al. (2023a) leverages GPT-4 to generate various types of instructions, including dialogues, detailed descriptions, and complex reasoning, based on textual descriptions of images. Building on this, Li et al. (2024b) further expands the data scale, leading to performance improvements. Tong et al. (2024) enhances model performance by increasing the dataset size and adjusting the data type ratios, while Li et al. (2024a) introduces a "high-quality knowledge learning" phase to further enrich the model’s knowledge base. In addition, several works explore using closed-source commercial models to generate synthetic instruction data, such as generating captions with GPT-4o or GPT-4v models Chen et al. (2023, 2024a) or OCR data Carter (2024), as well as conversation data Wang et al. (2023a). Despite these advancements, existing open-source data and instruction datasets remain insufficient to support models in achieving optimal performance. Models trained solely on open-source data still significantly lag behind SOTA closed-source models or open-source models trained on proprietary data. The limitations in both the quantity and quality of open-source data are key factors constraining model performance.
@@
To further enhance the performance of open-source models, this work explores improving model effectiveness by expanding the scale of instruction data and increasing the diversity of instruction types. We have extensively collected existing open-source multimodal instruction data, constructing a dataset of approximately 40 million samples, and applied rigorous quality filtering and deduplication processes. The model trained on this dataset demonstrated excellent performance, achieving a very high level of accuracy. Building on this, we propose a multimodal instruction synthesis method based on open-source VLM models. By providing highly detailed annotations for images and generating diverse questions for each image to ensure comprehensive coverage of the information, we can produce higher-quality instruction data, further improving the model’s ability to understand and follow instructions. Ultimately, we successfully trained a 2-billion-parameter VLM model based on open-source data and synthetic data generated by open-source models, achieving SOTA performance comparable to SOTA open-source models of similar scale.
The key contributions of this research include:
@@
  • We collected, organized, and open-sourced a large-scale multimodal instruction dataset, Infinity-MM, consisting of tens of millions of samples. Through quality filtering and deduplication, we ensured the dataset’s high quality and diversity.
  • We proposed a synthetic data generation method based on open-source models and a labeling system, capable of producing high-quality instruction data and effectively expanding the scale of instruction datasets.
@@
  • Based on Infinity-MM, we successfully trained a 2-billion-parameter VLM model, Aquila-VL-2B, achieving state-of-the-art performance among models of the same scale.
@@ 2 Related Work
Vision-Language Model VLMs can be categorized into three types based on their capabilities. The first type focuses on understanding multimodal information, such as videos and images Radford et al. (2021); Alayrac et al. (2022); Liu et al. (2024b); Li et al. (2023); Diao et al. (2024). These models typically take multimodal data as input and produce natural language output, characterized by their ability to integrate and process information from different modalities in a unified manner. The second type emphasizes visual generation, primarily aimed at producing high-resolution images and videos Shi et al. (2020); Peebles and Xie (2023); Ramesh et al. (2021); Ding et al. (2021). The third type combines both visual understanding and generation capabilities Sun et al. (2024b, a); Wang et al. (2024b); Zhou et al. (2024); Xie et al. (2024). In this work, we focus on enhancing the model’s ability to comprehend multimodal information.
Multi-modal Instruction Data Currently, a considerable amount of research has explored leveraging closed-source commercial models (mainly the GPT-4 series) to generate synthetic instruction data. The first category of work primarily utilizes GPT-4o or GPT-4v to generate specific types of data, such as captions Chen et al. (2023, 2024a), OCR Carter (2024) and conversations Wang et al. (2023a). Another category of work attempts to generate more complex dialogue or other types of instructions. For example, Liu et al. (2023a) uses GPT-4 to generate various types of instructions based on textual descriptions of images. Wang et al. (2023a) directly uses GPT-4V to generate instructions from images, though it generates only one instruction per image except for the text description for the image. In this work, we focus on how to leverage open-source models to generate high-quality multimodal instruction data.
3 Data
@@
Image-Cpation Data 10M Caption Data 10M
General Visual Instruction Data 24.4M
General Data 7.1M
General OCR Data 2.6M
@@
Text Instruct Data 7.6M
Selective Visual Instruction Data 6M
LLaVA-onevision Data 3.5M
Infinity-Instruct(subjective part) 1.3M
Docmatix Data 1.2M
@@
Data Generated by GPT4 1.7M
Synthetic Data 0.8M
Specific Task Data 0.4M
Infinity-Preference Data 0.1M
Table 1: The quantity and composition of the training data.
First, we extensively collect existing open-source multimodal datasets and categorize them based on task and quality. Subsequently, we will introduce the process of synthesizing data. Finally, we performed a unified deduplication and filtering of all collected data. The next three subsections will each address specific aspects in detail.
3.1 Categories of Multimodal Datasets
@@
We systematically gathered available open-source multimodal datasets and categorized them. These datasets were classified into four categories, as outlined in Table 1.
  • Image-Caption Data We collected the Image-Caption dataset generated by Emu2 (Sun et al., 2024a). Caption generation is a relatively fundamental task, making it well-suited for the initial training of large multimodal models.
  • General Visual Instruction Data We collected various general task data encompassing OCR, mathematical reasoning, chart comprehension, and other tasks. Training large multimodal models with this data equips them with the fundamental capabilities to tackle multimodal tasks effectively.
  • Selective Visual Instruction Data The sources of this data are Llava-OneVision Li et al. (2024a), Docmatix Laurençon et al. (2024) and the subjective components of Infinity-Instruct BAAI (2024b). Verifications on these data have shown that the quality of this data is superior to that of general task instruction data.
  • GPT4 & Synthetic Data This part mainly includes data generated by GPT-4 and the synthetic instruction data introduced in Section 3.2, along with a small amount of data specifically tailored for targeted tasks. Experimental results indicate that training with datasets containing synthetic data can further enhance model performance.
3.2 Synthetic Data Generation
@@
In this study, we propose a multimodal instruction data synthesis method based on open-sourced VLMs. Our goal is to ensure that the generated instructions are closely aligned with the content of the images, while maintaining diversity in instruction types and ensuring the accuracy of instruction responses. The overall process of the method is shown in Figure 1. The images of the synthetic data are extracted from the instruction dataset synthesized using the GPT-4 series models, which is of high quality. However, due to budget constraints, the scope and quantity of the synthetic data are limited. Therefore, we aim to leverage open-source models to synthesize more high-quality data, combining it with the original data to further enhance model performance.
@@ 3.2.1 Image and Instruction Tagging System
We first utilize the RAM++ model Huang et al. (2023) to automatically annotate images by extracting key information such as objects, actions, and scenes. These tags form the semantic foundation of the images, providing a critical basis for subsequent instruction generation. The RAM++ model demonstrates excellent performance when processing large-scale image datasets, accurately capturing essential details in multimodal scenes. This lays a solid foundation for generating precise and contextually relevant multimodal instructions.
To systematize the instruction generation process, we designed a three-level instruction tagging system that covers different types of instructions. Following Liu et al. (2023b), the first-level tags of the instruction tagging system are divided into six categories, which are:
@@
The middle level further refines task characteristics, while the bottom level provides a detailed classification based on specific task requirements. We employed a commercial closed-source model to extend and enhance this system, ensuring its comprehensiveness and rationality. The complete tagging system can be found in Appendix D.
Figure 1: Illustration of synthetic data generation method.
@@ 3.2.2 Question Generation
We randomly selected a portion of the open-source data we collected as seed data and annotated both the images and instructions using the method described in the previous section. We then established a set of mapping rules by analyzing the correlations between image tags and instruction tags. Specifically, we calculated the TF-IDF values for the image tags corresponding to each instruction type and ranked the results. Higher TF-IDF values indicate that images with those tags are more suitable for generating that particular type of instruction. Using these rules, we can automatically determine the appropriate instruction type to generate when processing new images. This approach significantly enhances the alignment between generated instructions and image content.
During question generation, we input both the images and the target instruction type into the VLM model, prompting the model to generate questions based on the image. Additionally, we randomly select two examples from the seed data to input into the model alongside the image for reference, enabling few-shot generation. For the questions generated by the VLM, we further input both the image and the question back into the question VLM to evaluate the relevance of the question to the image, filtering out lower-quality questions.
@@ 3.2.3 Answer Generation
Figure 2: The distribution of instruction types of the synthetic data.
After generating the questions, we proceeded to generate the corresponding instruction answers. The goal at this stage was not only to ensure the accuracy of the generated answers but also to account for the diversity of different instruction types. To achieve this, we introduced various prompts to increase answer diversity. Specifically, we employed three different types of prompts: one instructed the model to provide short answers using single words or phrases; another prompted the model to first generate a simple explanation before giving the answer; and the third prompted the model to provide a detailed explanation followed by the answer. We then input the image, question, and generated answer into the VLM model to filter out instructions and answers that did not align with the image content or task.
Finally, we obtained approximately 10M question-answer pairs. To further ensure the quality of the generated data, we input the images, questions, and answers into the Qwen2-VL-2B model to compute the data loss and filtered out about 3M samples. We combined multiple QA pairs corresponding to the same image into multi-turn instruction data, resulting in approximately 800K training instructions. The distribution of instruction types in the final synthetic data is shown in Figure 2.

### v2

@@ Abstract
Recently, Vision-Language Models (VLMs) have achieved remarkable progress in multimodal tasks, and multimodal instruction data serves as the foundation for enhancing VLM capabilities. Despite the availability of several open-source multimodal datasets, limitations in the scale and quality of open-source instruction data hinder the performance of VLMs trained on these datasets, leading to a significant gap compared to models trained on closed-source data. To address this challenge, we introduce Infinity-MM, a large-scale multimodal instruction dataset. We collected the available multimodal instruction datasets and performed unified preprocessing, resulting in a dataset with over 40 million samples that ensures diversity and accuracy. Furthermore, to enable large-scale expansion of instruction data and support the continuous acquisition of high-quality data, we propose a synthetic instruction generation method based on a tagging system and open-source VLMs. By establishing correspondences between different types of images and associated instruction types, this method can provide essential guidance during data synthesis. Leveraging this high-quality data, we have trained a 2-billion-parameter Vision-Language Model, Aquila-VL-2B, which achieves state-of-the-art (SOTA) performance among models of similar scale. The data is available at: https://huggingface.co/datasets/BAAI/Infinity-MM.
@@ 1 Introduction
Recently, Vision-Language Models (VLMs) [35, 41, 16, 80, 7, 64, 68, 51, 70, 63, 14, 33] have made significant progresses and found important applications across numerous fields, drawing increasing attention. With the development in foundational language models, multimodal architectures, multimodal training data, and evaluation benchmarks, the capabilities of multimodal models have greatly improved. Among these elements, multimodal instruction data forms the foundation of a multimodal model’s capabilities. Expanding training data, enhancing data quality, and increasing data diversity play crucial roles in advancing model performance [41, 40, 60, 33, 34].
Figure 1: Average score of different VLMs on benchmarks. The Aquila-VL-2B model, trained with Infinity-$M$, not only outperforms models trained on other open-source datasets (OneVision-SI) but also surpasses models trained on closed-source datasets.
@@
Many works have focused on exploring more effective ways to generate and utilize training data. For instance, Liu et al. [40] leverages GPT-4 to generate various types of instructions based on textual descriptions of images. Building on this, Li et al. [34] further expands the data scale, leading to performance improvements. Tong et al. [60] and Li et al. [33] enhance model performance by increasing the dataset size and adjusting the data type ratios. Besides, several works explore using GPT-4 series models to generate synthetic instruction data, such as captions [12, 10, 36] , OCR data [9], math questions [55], and conversation data [61, 10, 38, 71]. Despite these advancements, existing open-source instruction datasets remain insufficient to support models in achieving best performance. Models trained solely on open-source data [60, 33] still significantly lag behind SOTA closed-source models [51, 4] or open-source models trained on closed-source data [63, 14, 70]. Compared to closed-source data, existing open-source data is significantly limited in scale, restricting the foundational capability enhancement of models [47]. Besides, current methods for acquiring high-quality open-source data often rely on commercial closed-source models, such as GPT-4v and GPT-4o, which can be costly, making large-scale data generation a challenge. Furthermore, the synthesized instruction types are typically confined to predefined categories or are generated freely by the model without sufficient guidance on instruction types, leading to limitations in correctness and diversity. The limitations in both the quantity and quality of open-source data are key factors constraining open model performance.
Table 1: The comparision between Infinity-MM and other multimodal instruction datasets. ”Size” denotes the overall volume of data in the dataset, ”Data Synthesis Method” denotes to the specific approach used to generate synthetic data within the dataset, and ”Data Composition” denotes the types of instruction data included in the dataset. Compared to existing open-source datasets, Infinity-MM has a significant advantage in data scale. Besides, this is the first time to use an open-source VLM for large-scale, high-quality instruction data synthesis.
Datasets Size Data Synthesis Methods Data Composition
from GPT4(o/v) from Open VLM General Instruction OCR Doc/Chart/Screen Math/Reasoning Text Instruction
LVIS-Instruct[25] 223K
Sharegpt4[12] 3.2M
MMC-INST[38] 500K
ALLaVA[10] 1.7M
MathV360K[55] 360K
DocStruct4M[66] 4M
DocDownstream [71] 700K
Llava-1.5[40] 660K
Cambrain-1[60] 10M
Llava-OneVision[33] 7.2M
Infinity-$M$ 44.8M
@@
To enhance the performance of open-source models, this work explores improving data effectiveness by expanding the scale of instruction data and increasing the diversity of instruction types. We have extensively collected existing open-source multimodal instruction data, constructing a dataset of over 40 million samples, and applied rigorous quality filtering and deduplication processes. Besides, we propose a data synthesis method based on a tagging system. We establish tagging systems for images and instructions, modeling the relationship between image and instruction types through this system. This provides essential guidance for data synthesis, ensuring both the accuracy and diversity of the synthesized data. We also employed open-source VLMs for large-scale, high-quality data synthesis, providing detailed and diverse annotations for images to cover all the information within them as comprehensively as possible, thereby enhancing the model’s performance. Ultimately, we successfully trained a 2-billion-parameter model Aquila-VL-2B based on the proposed dataset Infinity-MM, which not only outperforms models trained on other open-source datasets but also surpasses models trained on closed-source datasets, as shown in Figure 1. The key contributions are summarized as follows:
@@
  • We collected and organized large-scale multimodal instruction dataset, Infinity-MM, consisting of tens of millions of samples. Through quality filtering and deduplication, we ensured the dataset’s high quality and diversity.
  • We propose a data synthesis method based on a tagging system, which provides essential guidance for data synthesis. We employed open-source VLMs to perform large-scale, high-quality instruction data synthesis.
@@
  • Based on Infinity-MM, we have successfully trained a 2-billion-parameter VLM model, Aquila-VL-2B, achieving SOTA performance among models of the same scale.
@@ 2 Related Work
Multi-modal Instruction Data Currently, the construction of multimodal instruction datasets primarily employs three methods: manual annotation, model synthesis, and open-source data collection. The first method is manual annotation. Das et al. [17] involves two annotators per image engaging in up to 10 rounds of question-answer interactions, generating multi-turn dialogues associated with each image. Agrawal et al. [2] employs one set of workers to create up to three QA pairs based on given images and text, followed by verification by a second set of workers to ensure accuracy. Havard et al. [26] combines manual annotation and OCR techniques to build a dataset with detailed text detection and recognition annotations. Mathew et al. [49] adopts a three-stage manual annotation process to create a visual question-answering dataset tailored for document images, featuring diverse question types and a wide range of document images. The second kind of methods utilize VLM model to synthesize specific types of data, such as captions [12, 10] , OCR data [9], math questions [55], chart description [71, 31], and conversation data [61, 10, 40]. The third method is open-source data collection. Liu et al. [38] collected 210K chart-title pairs by downloading academic articles from 2010 to 2020. Ye et al. [71] aggregates multiple publicly available document image datasets into a unified, diverse document dataset. Tong et al. [60] and Li et al. [33] both compile a variety of available open-source instruction datasets to create a collection encompassing multiple data types. Our proposed Infinity-MM dataset combines model synthesis with open-source data collection. As shown in Table 1, compared to existing datasets, Infinity-MM includes various types of data and has a substantial advantage in data scale.
Vision-Language Model VLMs can be categorized into three types based on their capabilities. The first type focuses on understanding multimodal information, such as videos and images [53, 3, 41, 35, 18]. These models typically take multimodal data as input and produce natural language output, characterized by their ability to integrate and process information from different modalities in a unified manner. The second type emphasizes visual generation, primarily aimed at producing high-resolution images and videos [56, 52, 54, 19]. The third type combines both visual understanding and generation capabilities [59, 58, 65, 79, 69]. In this work, we focus on enhancing the model’s ability to comprehend multimodal information.
3 The proposed Infinity-MM Dataset
Table 2: The data quantity and types contained in Infinity-$M$. We segmented the dataset based on data type and quality, allowing the model to perform targeted training at different stages.
@@ 3 The proposed Infinity-MM Dataset
Image Caption Data 10M General Instruction Data 10M
Visual Instruction Data Comprehensive Data 25.8M
General Instruction Data 7.1M
OCR Data 2.6M
@@
Text Instruction Data 9M
Selective Data 6M
General Instruction Data 1.3M
OCR Data 0.3M
Doc/Chart/Screen Data 1.9M
Math/Reasoning Data 0.7M
Text Instruction Data 1.8M
@@
General Instruction Data 1M
OCR Data 0.5M
Doc/Chart/Screen Data 0.1M
Math/Reasoning Data 0.3M
Text Instruction Data 0.3M
Our Synthetic Data 0.8M
Figure 2: Illustration data synthesis method. We propose a tagging system to model the correspondence between image types and instructions, providing essential guidance for data synthesis to ensure both accuracy and diversity in instruction generation. Besides, we introduce the use of open-source VLM for large-scale, high-quality instruction data synthesis.
In this section, we provide a detailed explanation of the process to construct the Infinity-MM dataset. First, we extensively collect available open-source multimodal datasets of various tasks and types and categorize them based on task and quality. Then, we propose a data synthesis method based on a tagging system. We establish tagging systems for both images and instructions within the multimodal instruction dataset. By modeling the relationship between image and instruction tags, we provide essential guidance during instruction synthesis, ensuring that the generated instructions align with the content of the images while also increasing diversity of the synthesized instructions. Finally, we standardize the formats of data from different sources and apply deduplication and quality filtering to the dataset.
3.1 Collection of Datasets
@@
We systematically gather available open-source multimodal datasets and categorize them. These datasets are classified into four categories, as outlined in Table 2. The specific sources of the data are given in Appendix B.
  • Image-Caption Data We collect the Image-Caption dataset generated by Emu2 [58], which performs well in image caption task. The image-caption data has detailed descriptions of image content, making this it well-suited for vision-language alignment training in VLMs.
  • Comprehensive Visual Instruction Data This part of the data is collected from open-source datasets and includes various types of data. The data distribution has not been specifically adjusted, and its effectiveness has not been rigorously validated. Therefore, this part of data is suitable for enhancing the foundational capabilities of VLMs.
  • Selective Visual Instruction Data This part of the data is also sourced from open-source datasets, with an adjusted data distribution that increases the proportion of tasks involving mathematical reasoning, chart analysis, and complex instructions—areas where VLMs generally perform poorly. The quality and effectiveness of these data have been validated, making them suitable for enhancing the instruction-following ability of VLMs.
  • GPT4 & Synthetic Visual Instruction Data We collect instruction data synthesized using GPT-4 series models. This data generally exhibits high quality, but the task types are relatively homogeneous, and the data volume is limited. We randomly sample a subset of images from this data and synthesized instructions based on these images with the method introduced in Section 3.2.
3.2 Data Synthesis
@@
In this section, we introduce our multimodal instruction data synthesis method. We aim that the generated instructions are closely aligned with the content of the images while maintaining diversity in instruction types and ensuring the accuracy of instruction responses. The overall process of the method is shown in Figure 2. The images of the synthetic data are extracted from the instruction dataset synthesized using the GPT-4 series models, which are of high quality. However, the instruction type and data quantity of the original data are limited. Therefore, we aim to leverage open-source models to synthesize more high-quality data, combining it with the original data to further enhance model performance.
@@ 3.2.1 Image and Instruction Tagging System
We randomly select a portion of the open-source data as seed data. Then we utilize the RAM++ model [29] to automatically annotate images by extracting key information such as objects, actions, and scenes. These tags form the semantic foundation of the images, providing a critical basis for subsequent instruction generation. The RAM++ model demonstrates excellent performance when processing large-scale image datasets, accurately capturing essential details in multimodal scenes. This lays a solid foundation for generating precise and contextually relevant multimodal instructions.
To systematize the instruction generation process, we design a three-level instruction tagging system that covers different types of instructions. Following Liu et al. [42], the first-level tags of the instruction tagging system are divided into six categories, which are:
@@
We extend the tagging system based on the first-level tags with the GPT-4 model. The middle level further refines task characteristics, while the bottom level provides a detailed classification based on specific task requirements. After constructing the initial tagging system, we use the GPT-4o model to annotate the instructions in the seed data. During the annotation process, we refined the tagging system by adding or removing tags. The resulting tagging system covers 199 sub-tasks, ensuring its comprehensiveness and rationality. The complete tagging system can be found in Appendix C.
Figure 3: This is an example of our synthesized data. Compared to the original data on the left, our synthesized data on the right offers greater diversity in instructions (Q&A 1-3) and covers a broader range of image content (Q&A 4,5).
3.2.2 Image-Instruction Mapping
We annotate both the images and instructions in the seed data with the method in the previous section. Following this, we establish the mapping relationships between image tags and instruction tags. Specifically, we begin by counting the frequency of image tags associated with each instruction tag in the seed data. Then, treating each instruction tag as a distinct unit, we calculate the TF-IDF values for the image tags within these units. A higher TF-IDF value for an image tag within a given instruction type indicates that images with this tag are more suitable for generating that specific type of instruction. Leveraging these TF-IDF-based mappings allows us to automatically determine the most appropriate instruction type to generate for new images. This approach significantly improves the alignment between the generated instructions and the actual image content.
@@ 3.2.3 Question Generation
After establishing the mapping relationships between images and instruction types, we proceed with instruction synthesis. Balancing data synthesis quality and efficiency, we employ the MiniCPM-V2.6 model [70] in this work. First, we need to generate appropriate questions for each image. For each candidate image, we identify the suitable instruction types. Then, we input both the image and the target instruction type into the VLM model, prompting it to generate questions based on these information. Besides, we select two data examples of the same instruction type from our annotated seed data and input their images and questions into the VLM for reference, enabling few-shot generation. For the questions generated by the VLM, we further re-input both the image and question into the VLM to evaluate the relevance of each question to the image, filtering out lower-quality questions.
@@ 3.2.4 Answer Generation
Figure 4: The distribution of first-level tags for synthetic instruction data. The tagging system enables more effective targeted synthesis and analysis of instruction data.
After generating the questions, we proceed to generate the corresponding instruction answers. We ensure not only the accuracy of the answers but also the diversity of their formats. To achieve this, we employ different prompts during the answer generation process. Specifically, we employ three different types of prompts: one instructed the model to provide short answers using single words or phrases; another prompted the model to first generate a simple explanation before giving the answer; and the third prompted the model to provide a detailed explanation followed by the answer.
We re-input the image, question, and answer into the VLM model, allowing it to score the instruction quality on a scale of 1 to 10, where a higher score indicates better quality. Instructions with scores below 8 were filtered out. Additionally, we input the synthesized instruction data into the Qwen2-VL-7B model [63] model to compute the loss, filtering out data with excessively high loss values. Finally, we select approximately 3 million filtered data points for training, prioritizing data related to reasoning and document analysis, as these tasks were relatively scarce in the previous training phases. We combine multiple QA pairs corresponding to the same image into multi-turn instruction data, resulting in approximately 800K training instructions. Figure 3 presents an example of our synthesized data. The distribution of instruction types in the final synthetic data is shown in Figure 4.